OpenStudy (chihiroasleaf):

Finding the value of limit \[\Large \lim_{x \to 3} \frac{\sqrt{x+4}-\sqrt{2x+1}}{\sqrt{3x+1}-\sqrt{x+7}}\]

OpenStudy (zehanz):

I would use l'Hôpital's Rule in this case.

OpenStudy (chihiroasleaf):

other than l'Hôpital's Rule?

OpenStudy (zehanz):

I started a calculation, but got stuck :( Therefore, I went to the emergency exit: l'Hôpital :)

OpenStudy (chihiroasleaf):

I've tried to multiply by the conjugates of the numerator and got stuck, then trying to multiply by the conjugates of the denumerator and got stuck too... am I allowed to multiply twice? by the conjugates of the numerator and denumerator?

OpenStudy (zehanz):

As long as you multiply num. and den. with the same number, anything goes. I also tried ot, but didn't get any further. That's why I came up with l'Hôpital. That always feels a bit like cheating, though...

OpenStudy (chihiroasleaf):

it seems doesn't work...

OpenStudy (experimentx):

multiply both by conjugate ... cancel (x-3) from both of them.

OpenStudy (zehanz):

That works! \[\frac{ (\sqrt{x+4}-\sqrt{2x+1})(\sqrt{x+4}+\sqrt{2x+1})(\sqrt{3x+1}+\sqrt{x+7}) }{ (\sqrt{3x+1}-\sqrt{x+7})(\sqrt{3x+1}+\sqrt{x+7})(\sqrt{x+4}+\sqrt{2x+1} ) }=\]\[\frac{ (x+4-(2x+1))(\sqrt{3x+1}+\sqrt{x+7}) }{ (3x+1-(x+7))(\sqrt{x+4}+\sqrt{2x+1} ) }=\]\[\frac{ -(x-3)(\sqrt{3x+1}+\sqrt{x+7}) }{ 2(x-3)(\sqrt{x+4}+\sqrt{2x+1} ) }=\]\[-\frac{ 1 }{ 2 }\cdot \frac{ (\sqrt{3x+1}+\sqrt{x+7}) }{ (\sqrt{x+4}+\sqrt{2x+1} ) }\]Now let x be 3, to see what the limit will be: \[-\frac{ 1 }{ 2 }\cdot \frac{ 2\sqrt{10} }{ 2\sqrt{7} }=-\frac{ 1 }{14 }\sqrt{70}\]

OpenStudy (chihiroasleaf):

ahh..., there will be cancellation... thanks a lot... :)